# Structured state-space models are deep Wiener models

### v1

@@ Abstract
The goal of this paper is to provide a system identification-friendly introduction to the Structured State-space Models (SSMs). These models have become recently popular in the machine learning community since, owing to their parallelizability, they can be efficiently and scalably trained to tackle extremely-long sequence classification and regression problems. Interestingly, SSMs appear as an effective way to learn deep Wiener models, which allows to reframe SSMs as an extension of a model class commonly used in system identification. In order to stimulate a fruitful exchange of ideas between the machine learning and system identification communities, we deem it useful to summarize the recent contributions on the topic in a structured and accessible form. At last, we highlight future research directions for which this community could provide impactful contributions.
@@
angle=0, hpos=0.5vpos=0.97fontsize=0.012color=[gray]0.2, text= This work is an extended version of a manuscript submitted to IFAC for possible publication. ,
@@ 1 Introduction
Recent years have been characterized by a remarkable research interest towards deep learning tools for data-driven control. An interesting use case is that of nonlinear system identification (Schoukens and Ljung, 2019). Over the years, a variety of (increasingly complex) Neural Network (NN) architectures have indeed been proposed for nonlinear system identification, ranging from gated Recurrent NNs (RNNs) (Bonassi et al., 2022) to Transformer NNs (TNNs) (Sun and Wei, 2022). These models have proven to work well in many challenging system identification problems, enabling the synthesis of accurate model predictive control laws, see e.g. Lanzetti et al. (2019); Terzi et al. (2020).
@@
Despite their modeling power, these architectures are known to suffer from computational efficiency problems during training. On the one hand, RNNs are inherently sequential models, meaning that they have to be iteratively unrolled over the time axis $T$ of training sequences (Bianchi et al., 2017). On the other hand, TNNs are plagued by a quadratic scaling issue ($\mathcal{O}(T^{2})$), thus calling for tailored, more efficient architectures like Longformers, which are however still untapped in the nonlinear system identification realm. Unfortunately, these problems have limited the use of RNNs and TNNs for long-term nonlinear system identification to those cases where enough computational budget is available.
@@
What is more, both RNNs and TNNs are still fairly unexplored when it comes to their system-theoretical properties, such as stability, which are however essential for utilizing these models for control design purposes (Bonassi et al., 2024). Most of the research in this area has been carried out for specific RNN models, see Miller and Hardt (2019); Bonassi et al. (2022), and mainly involve stability promotion via regularization.
@@
To address these problems, Gu et al. (2021) proposed a Structured State-space Model (SSM) architecture named S4, which consists of multiple layers composed by LTI discrete-time systems followed by a nonlinear function. The term “structured” stems from the fact that this LTI system is given a specific structure to improve the architecture’s modeling performances while also reducing the computational cost during training (Yu et al., 2018). Nonlinear state-space models are not new, see Marconato et al. (2013), yet their adoption has been hampered by their crucial reliance on the model structure and on the initialization method of learnable parameters. The contribution of the S4 approach towards SSMs has therefore been that of providing (i) a novel, intrinsically stable, parametrization of the LTI system obtained by discretizing a continuous-time Diagonal Plus-Low Rank (DPLR) system, (ii) a new strategy towards the parameters’ initialization problem, (iii) a computationally efficient approach to simulate (and train) these models over extremely long sequences, and (iv) an empirical proof of the state-of-the-art performances of these models in long-term sequence learning problems.
@@
Motivated by these appealing features, many works have built on the S4 architecture. For example, Gupta et al. (2022) and Gu et al. (2022) have explored the benefits entailed by stricter SSM structures, namely the parametrization via diagonal continuous-time systems (S4D), and by simpler initialization strategies. Smith et al. (2022) have explored a novel, and somewhat more computationally efficient, simulation method for diagonal continuous-time parametrizations, named S5. Orvieto et al. (2023) recently investigated the parametrization of the LTI subsystems directly in the discrete time domain, resulting in the Linear Recurrent Unit (LRU) architecture.
@@
Despite the appealing results achieved by SSMs in the long-range arena benchmarks sequence classification problems, their use for nonlinear system identification is still unexplored. With this paper we want to change that by making the following contributions. First of all we show that it is possible to interpret SSMs as deep Wiener models, i.e. model structures where several Wiener models are interconnected in series. An interesting note here is that even though the Wiener models have been extremely popular within system identification — see e.g. Schoukens and Tiels (2017) and references therein — their structure has been limited to the “single-layer”. Our second contribution is to dissect the recent developments on SSMs and explain them in terms of their structure and parameterization, and to clearly separate this from their initialization, simulation, and training strategies. The presentation in the paper is also done using the language commonly used in the system identification community in order to speed up the use of these tools within this area.
Notation
The imaginary unit is denoted by $i=\sqrt{-1}$. Given a vector $v$, we denote by $v^{\prime}$ its real transpose. For a time time-dependent vector, the discrete-time index $k$ is reported as a subscript, e.g., $v_{k}$. Moreover, we denote by $v_{a:b}$ (where $a\leq b$) the sequence
$v_{a:b}=\big{(}v_{a},v_{a+1},...,v_{b}\big{)}.$
@@
For this sequence, we indicate by $\operatorname{cat}(v_{a:b})$ the concatenation of its elements, i.e., $\operatorname{cat}(v_{a:b})=[v_{a}^{\prime},...,v_{b}^{\prime}]^{\prime}$. Given a (complex-valued) matrix $A$, we let $\operatorname{conj}(A)$ be its element-wise complex conjugate and $A^{*}$ be its Hermitian transpose. Diagonal matrices may be defined via the $\operatorname{diag}$ operator, as $A=\operatorname{diag}(a_{1},...,a_{n})$.
@@ 2 Structured State-space Models
Consider the model depicted in Figure 1, which consists of $L$ Wiener systems interconnected in series. Each of these layers is here referred to as Structured State-space Layer (SSL). Their interconnection results in a SSM, which can be interpreted as a specific configuration of deep Wiener system. We let the generic $\ell$-th SSL ($\ell\in\{1,...,L\}$) be represented by a discrete-time state-space model
$\mathcal{S}_{\ell}:\,\begin{dcases}x_{k+1}=Ax_{k}+Bu_{k},\\ \eta_{k}=Cx_{k}+Du_{k},\\ y_{k}=\sigma(\eta_{k})+Fu_{k},\end{dcases}$ (1)
where, for compactness, the layer index is omitted. System (1) is characterized by the input vector $u\in\mathbb{R}^{n_{u}}$, the intermediate vector $\eta\in\mathbb{R}^{n_{y}}$, the output vector $y\in\mathbb{R}^{n_{y}}$, and the complex-valued state vector $x\in\mathbb{C}^{n_{x}}$. This system is parametrized by the matrices $\{A,B,C,D,F\}$. The nonlinear output transformation $\sigma(\cdot)$ can be any nonlinear activation function, such as the $\tanh$, ELU, or Swish, see Ramachandran et al. (2017). In what follows, we aim to provide an overview of the possible structure, parametrization, initialization, and simulation strategies for this SSL.
@@
We start by observing that, because both the input and output vector of the SSL are real valued, the system matrices $A$, $B$, and $C$ of (1) might be re-written as
@@
where $\tilde{A}\in\mathbb{C}^{n_{\lambda}\times n_{\lambda}}$, with $n_{\lambda}=\frac{n_{x}}{2}$, $\tilde{B}\in\mathbb{C}^{n_{\lambda}\times n_{u}}$, and $\tilde{C}\in\mathbb{C}^{n_{y}\times n_{\lambda}}$, while $D\in\mathbb{R}^{n_{y}\times n_{u}}$. Matrix $F\in\mathbb{R}^{n_{y}\times n_{u}}$ might be interpreted as a skip connection, and it is often fixed to $F=I$ whenever $n_{u}=n_{y}$. Structure (2) ensures that the eigenvalues come in complex-conjugate pairs.
Even under (2), both in the linear case (Yu et al., 2018) and in the nonlinear case (Marconato et al., 2013) it is well known that state-space models call for additional structure to reduce the number of learnable parameters and hence achieve improved modeling performances and computational efficiency. The initialization of such learnable parameters is also known to be paramount to achieve satisfactory performances. In the following, the parametrization and initialization strategies proposed in recent literature are discussed.
@@ 3.1 Discrete-time diagonal parametrization
Note that this parametrization structurally guarantees the asymptotic stability of the SSL, since $\lvert\lambda_{j}\lvert<1$. Another relevant design choice advocated by Orvieto et al. (2023) is that of reparametrizing $\tilde{B}$ as
$\tilde{B}=\operatorname{diag}(\gamma_{1},...,\gamma_{n_{\lambda}})\tilde{\tilde{B}},$ (4a)
@@
This ensures that the rows of $\tilde{B}$ are normalized, so that white noise inputs yield state trajectories with the same energy content as of the input. Therefore, the set of learnable parameters of this SSL parametrization reads
$\Theta=\Big{\{}\{\mu_{j},\theta_{j}\}_{j\in\{1,...,n_{\lambda}\}},\tilde{\tilde{B}},\tilde{C},D,F\Big{\}}.$ (5)
@@ 3.2 Initialization strategy
where $\mathfrak{U}[a,b]$ denotes the uniform distribution with support $[a,b]$. The complex matrices $\tilde{\tilde{B}}$ and $\tilde{C}$, and the real matrices $D$ and $F$ [^1] We remind that if $n_{u}=n_{y}$ the skip connection $F$ is usually fixed to the identity matrix. can be initialized with the Xavier initialization method (Kumar, 2017). That is, they are sampled from a Normal distribution whose variance is scaled by a factor proportional to the number of columns.
@@ 4 Continuous-time reparametrizations
Another viable approach is that of parametrizing the SSL via a discretized continuous-time system, as proposed by Gu et al. (2021) in the context of S4 models. To this end, we let the SSL be parametrized in the continuous-time domain as
@@
with $\tilde{A}_{c}\in\mathbb{C}^{n_{\lambda}\times n_{\lambda}}$, $\tilde{B}_{c}\in\mathbb{C}^{n_{\lambda}\times n_{u}}$, $\tilde{C}_{c}\in\mathbb{C}^{n_{y}\times n_{\lambda}}$. The term $\Gamma$ in (8) is a learnable timescale parameter, which can either be a real and positive scalar (Gu et al., 2021) or a real-valued, positive-definite diagonal matrix (Smith et al., 2022), i.e., $\Gamma=\operatorname{diag}(\gamma_{1},...,\gamma_{n_{\lambda}})$.
Under this parametrization, the discrete-time system matrices $\{{A},{B},{C},D\}$ can finally be expressed in terms of their continuous-time counterparts, $\{A_{c},B_{c},C_{c}\}$, by defining the sampling time $\tau$ and a discretization method, such as forward Euler, the bilinear transform, and Zero-Order Hold (ZOH). As shown in Section 3.1, additional structure can now be imposed on $\tilde{A}_{c}$ to mitigate the overparametrization, thus making the training procedure more scalable, and to enforce the structural stability of the SSL. The possible structures of $\tilde{A}_{c}$ are now discussed.
@@ 4.1 Continuous-time diagonal parametrization
Note that, because $\mathfrak{Re}(\lambda_{j})<0$, this parametrization structurally guarantees the asymptotic stability of the SSL. Overall, the set of learnable parameters is
@@
As a side note, it is worth pointing out that the eigenvalues can also be parametrized in terms of the modulus and phase instead, see (3b).
@@ 4.2 Continuous-time DPLR parametrization
where $\tilde{\Lambda}_{c}$ is a diagonal component defined as in (4), while $\tilde{P},\tilde{Q}\in\mathbb{C}^{n_{\lambda}\times n_{r}}$ yield a low rank component, with $n_{r}\ll n_{\lambda}$. More specifically, Gu et al. (2021) propose to adopt $n_{r}=1$ and $\tilde{P}=\tilde{Q}$. Note that this choice guarantees the asymptotic stability of the SSL parametrization, since the negative-definiteness of $\mathfrak{Re}(\tilde{\Lambda}_{c})$ implies that of $\mathfrak{Re}(\tilde{A}_{c})$. The set of learnable parameters, under the DPLR parametrization, is

### v2

@@ Abstract
The goal of this paper is to provide a system identification-friendly introduction to the Structured State-space Models (SSMs). These models have become recently popular in the machine learning community since, owing to their parallelizability, they can be efficiently and scalably trained to tackle extremely long sequence classification and regression problems. Interestingly, SSMs appear as an effective way to learn deep Wiener models, which allows us to reframe SSMs as an extension of a model class commonly used in system identification. To stimulate a fruitful exchange of ideas between the machine learning and system identification communities, we deem it useful to summarize the recent contributions on the topic in a structured and accessible form. At last, we highlight future research directions for which this community could provide impactful contributions.
@@
angle=0, hpos=0.5vpos=0.97fontsize=0.012color=[gray]0.2, text= ©2024 the authors. This work has been accepted to IFAC for publication under a Creative Commons Licence CC-BY-NC-ND. ,
@@ 1 Introduction
Recent years have been characterized by a remarkable research interest towards deep learning tools for data-driven control. An interesting use case is that of nonlinear system identification. Over the years, a variety of (increasingly complex) Neural Network (NN) architectures have indeed been proposed for nonlinear system identification, ranging from gated Recurrent NNs (RNNs) (Bonassi et al., 2022) to Transformer NNs (TNNs) (Sun and Wei, 2022). These models have proven to work well in many challenging identification problems, enabling the synthesis of accurate model predictive control laws, e.g. Lanzetti et al. (2019).
@@
Despite their modeling power, these architectures are known to suffer from computational efficiency problems at training. On the one hand, RNNs are inherently sequential models, implying they have to be iteratively unrolled over the time axis $T$ of training sequences (Bianchi et al., 2017). On the other hand, TNNs are plagued by a quadratic scaling issue ($\mathcal{O}(T^{2})$), thus calling for tailored architectures like Longformers, which are however still untapped in the nonlinear system identification realm. Unfortunately, these problems have limited the use of RNNs and TNNs for long-term nonlinear system identification to those cases where enough computational budget is available.
@@
What is more, both RNNs and TNNs are still fairly unexplored when it comes to their system-theoretical properties, such as Incremental Input-to-State Stability ($\delta$ISS) (Bonassi et al., 2022), which are however essential for utilizing these models for control design (Bonassi et al., 2024). Most of the research in this area is tailored to specific RNNs, see Miller and Hardt (2019); Bonassi et al. (2022), and mainly involves stability promotion via regularization.
@@
To address these problems, Gu et al. (2021) proposed a Structured State-space Model (SSM) architecture named S4, which consists of multiple layers composed by LTI discrete-time systems followed by a nonlinear function. The term “structured” stems from the fact that this LTI system is given a specific structure to improve the architecture’s modeling performances while also reducing the computational cost at training (Yu et al., 2018). Nonlinear state-space models are not new, see Marconato et al. (2013), yet their adoption has been hampered by their crucial reliance on the model structure and on the initialization method of learnable parameters. The contribution of the S4 approach towards SSMs has therefore been that of providing (i) a novel, intrinsically stable, parametrization of the LTI system obtained by discretizing a continuous-time Diagonal Plus-Low Rank (DPLR) system, (ii) a new strategy towards the parameters’ initialization problem, (iii) a computationally efficient approach to simulate (and train) these models over extremely long sequences, and (iv) an empirical proof of the state-of-the-art performances of these models in long-term sequence learning problems.
@@
Motivated by these appealing features, many works have continued to build on the S4 architecture. For example, Gupta et al. (2022) and Gu et al. (2022) have explored the benefits entailed by stricter SSM structures, namely the parametrization via diagonal continuous-time systems (S4D), and by simpler initialization strategies. Smith et al. (2022) have explored a novel, and somewhat more computationally efficient, simulation method for diagonal continuous-time parametrizations, named S5. Orvieto et al. (2023) recently investigated the parametrization of the LTI subsystems directly in the discrete time domain, resulting in the Linear Recurrent Unit (LRU) architecture.
1.0.1 Contribution
@@
Despite the appealing results achieved by SSMs in the long-range arena benchmarks sequence classification problems, their use for nonlinear system identification is still unexplored. With this paper, we want to change that by making the following contributions. First of all we show that it is possible to interpret SSMs as deep Wiener models, i.e. model structures where several Wiener models are interconnected in series. An interesting note here is that even though the Wiener models have been extremely popular within system identification — see e.g. Schoukens and Tiels (2017) and references therein — their structure has been limited to “single-layer” or parallel architectures (Wills and Ninness, 2012). Our second contribution is to dissect the recent developments on SSMs and explain them in terms of their structure and parameterization, and to clearly separate this from their initialization, simulation, and training strategies. The presentation in the paper is also done using the language commonly used in the system identification community in order to speed up the use of these tools within this area.
1.0.2 Notation
@@
The imaginary unit is denoted by $i=\sqrt{-1}$. Given a vector $v$, we denote by $v^{\prime}$ its real transpose. For a time-dependent vector, the discrete-time index $k$ is reported as a subscript, e.g., $v_{k}$. Moreover, we denote by $v_{a:b}$ (where $a\leq b$) the sequence $v_{a:b}=\big{(}v_{a},v_{a+1},...,v_{b}\big{)}$. For this sequence, we indicate by $\operatorname{cat}(v_{a:b})$ the concatenation of its elements, i.e. $\operatorname{cat}(v_{a:b})=[v_{a}^{\prime},...,v_{b}^{\prime}]^{\prime}$, and by $\operatorname{uncat}(\cdot)$ its inverse operation returning a sequence of vectors given their concatenation. Given a complex matrix $A$, we let $\operatorname{conj}(A)$ be its element-wise complex conjugate and $A^{*}$ be its Hermitian transpose. Diagonal matrices may be defined via the $\operatorname{diag}$ operator, as $A=\operatorname{diag}(a_{1},...,a_{n})$.
@@ 2 Structured State-space Models
Consider the model depicted in Figure 1, which consists of $L$ Wiener systems interconnected in series. Each of these layers is here referred to as Structured State-space Layer (SSL). Their interconnection results in an SSM, which can be interpreted as a specific configuration of a deep Wiener system. We let the generic $\ell$-th SSL ($\ell\in\{1,...,L\}$) be represented by a discrete-time state-space model
$\text{SSL}_{\ell}:\,\begin{dcases}x_{k+1}=Ax_{k}+Bu_{k},\\ \eta_{k}=Cx_{k}+Du_{k},\\ y_{k}=\sigma(\eta_{k})+Fu_{k},\end{dcases}$ (1)
where, for compactness, the layer index is omitted. System (1) is characterized by the input vector $u\in\mathbb{R}^{n_{u}}$, the intermediate vector $\eta\in\mathbb{R}^{n_{y}}$, the output vector $y\in\mathbb{R}^{n_{y}}$, and the complex-valued state vector $x\in\mathbb{C}^{n_{x}}$. The SSL is parametrized by the matrices $\{A,B,C,D,F\}$. The output transformation $\sigma(\cdot)$ can be any nonlinear, Lipschitz-continuous activation function, such as the $\tanh$, ELU, or Swish, see Ramachandran et al. (2017). In what follows, we aim to provide an overview of the possible structure, parametrization, initialization, and simulation strategies for this SSL.
@@
We start by observing that, because both the input and output vector of the SSL are real-valued, the system matrices $A$, $B$, and $C$ of (1) might be re-written as
@@
where $\tilde{A}\in\mathbb{C}^{n_{\lambda}\times n_{\lambda}}$, $\tilde{B}\in\mathbb{C}^{n_{\lambda}\times n_{u}}$, $\tilde{C}\in\mathbb{C}^{n_{y}\times n_{\lambda}}$, while $D\in\mathbb{R}^{n_{y}\times n_{u}}$, with $n_{\lambda}=\frac{n_{x}}{2}$. Matrix $F\in\mathbb{R}^{n_{y}\times n_{u}}$ can be interpreted as a skip connection, fixed to $I$ if $n_{u}=n_{y}$. Note that if the SSLs are parametrized to be structurally Schur stable, the stability of the SSM is guaranteed.
Proposition 2
If each layer $\ell\in\{1,...,L\}$ is parametrized by a Schur-stable matrix $\tilde{A}$, then the SSM is Incrementally Input-to-State Stable ($\delta$ISS) (Bonassi et al., 2022, Def. 4).
{pf}
Owing to the Schur stability of $\tilde{A}$ and Lipschitz continuity of $\sigma(\cdot)$, the SSL can be easily proven (globally) $\delta$ISS. The SSM is then the series of $\delta$ISS systems and is therefore $\delta$ISS (Angeli, 2002). $\hfill\qed$
While (2) ensures that the eigenvalues come in complex-conjugate pairs, both in the linear (Yu et al., 2018) and the nonlinear realm (Marconato et al., 2013) it is well known that state-space models call for additional structure to reduce the number of learnable parameters and hence achieve improved modeling performances and computational efficiency. Besides, the initialization of these parameters is also paramount to achieving satisfactory performances. In the following, the strategies proposed in recent literature to address these problems are discussed.
@@ 3.1 Discrete-time diagonal parametrization
Note that since $\lvert\lambda_{j}\lvert<1$ this parametrization guarantees that the SSM is structurally $\delta$ISS (Proposition 2). A relevant design choice advocated by Orvieto et al. (2023) is that of reparametrizing $\tilde{B}$ as
$\tilde{B}=\operatorname{diag}(\gamma_{1},...,\gamma_{n_{\lambda}}){\check{B}},$ (4a)
@@
This ensures that the rows of $\tilde{B}$ are normalized so that white noise inputs yield state trajectories with the same energy content as of the input. Therefore, the set of learnable parameters of this SSL parametrization reads
$\Theta=\Big{\{}\{\mu_{j},\theta_{j}\}_{j\in\{1,...,n_{\lambda}\}},\check{B},\tilde{C},D,F\Big{\}}.$ (5)
@@ 3.2 Initialization strategy
where $\mathfrak{U}[a,b]$ denotes the uniform distribution with support $[a,b]$. The complex matrices $\check{B}$ and $\tilde{C}$, and the real matrices $D$ and $F$ [^1] If $n_{u}=n_{y}$ the skip connection $F$ is usually fixed to the identity. can be initialized with the Xavier initialization method (Kumar, 2017). That is, they are sampled from a Normal distribution whose variance is scaled by a factor proportional to the number of columns.
@@ 4 Continuous-time reparametrizations
Another approach is that of parametrizing the SSL via a discretized continuous-time system, as proposed by Gu et al. (2021) in the context of S4 models. To this end, we let the SSL be parametrized as
@@
with $\tilde{A}_{c}\in\mathbb{C}^{n_{\lambda}\times n_{\lambda}}$, $\tilde{B}_{c}\in\mathbb{C}^{n_{\lambda}\times n_{u}}$, $\tilde{C}_{c}\in\mathbb{C}^{n_{y}\times n_{\lambda}}$. The factor $\Gamma$ in (8) is a learnable timescale parameter, which can either be a real and positive scalar (Gu et al., 2021) or a real-valued, positive-definite diagonal matrix (Smith et al., 2022), i.e., $\Gamma=\operatorname{diag}(\gamma_{1},...,\gamma_{n_{\lambda}})$.
Under this parametrization, the discrete-time system matrices $\{{A},{B},{C},D\}$ can finally be expressed in terms of their continuous-time counterparts, $\{A_{c},B_{c},C_{c}\}$, by defining the sampling time $\tau$ and a discretization method, such as forward Euler, the bilinear transform, or Zero-Order Hold (ZOH). As shown in Section 3.1, an additional structure can now be imposed on $\tilde{A}_{c}$ to mitigate the overparametrization, thus making the training procedure more scalable, and to enforce the structural stability of the SSL and hence, owing to Proposition 2, the stability of the SSM. The possible structures of $\tilde{A}_{c}$ are now discussed.
@@ 4.1 Continuous-time diagonal parametrization
Note that, because $\mathfrak{Re}(\lambda_{j})<0$, this parametrization structurally guarantees the asymptotic stability of the SSL, and hence the $\delta$ISS of the SSM (Proposition 2). Overall, the set of learnable parameters is
@@
As a side note, it is worth pointing out that the eigenvalues can also be parametrized in terms of the modulus and phase instead, as in (3b).
@@ 4.2 Continuous-time DPLR parametrization
where $\tilde{\Lambda}_{c}$ is a diagonal component defined as
$\tilde{\Lambda}_{c}=\operatorname{diag}(\lambda_{1},...,\lambda_{n_{\lambda}})$ (12a)
and $\tilde{P},\tilde{Q}\in\mathbb{C}^{n_{\lambda}\times n_{r}}$ yield a low rank component, with $n_{r}\ll n_{\lambda}$. In order to guarantee the structural stability of the SSL, and hence the $\delta$ISS of the SSM, one can take $\tilde{P}=\tilde{Q}$ (see Gu et al. (2021)) and
$\lambda_{j}=-\phi(\alpha_{j}^{\text{re}})+i\alpha_{j}^{\text{im}},$ (12b)
with $\phi(\cdot)$ being any strictly-positive function, applied element-wise, e.g. $\phi(\alpha)=\max(0,\alpha)+\varepsilon$, with $\varepsilon>0$. This ensures the negative definiteness of $\tilde{A}_{c}$.
The set of learnable parameters, under the DPLR parametrization, is